The balanced combustion reaction for C6H6 is

2C6H6(l)+15O2(g) --> 12CO2(g)+6H2O(l)+6542kJ

If 8.000 g C6H6 is burned and the heat produced from the burning is added to 5692 g of water at 21 C, what is the final temperature of the water?

final temperature: _________________ C
Answer #1

no of moles of C6H6   = W/G.M.Wt

                                      = 8/78   = 0.1026mole

2C6H6(l)+15O2(g) --> 12CO2(g)+6H2O(l)+6542kJ

\DeltaH   = -6542KJ

2 mole of C6H6 combustion to release energy is -6542KJ

0.1026 moles of C6H6 combustion to release energy = -6542*0.1026/2   = -335.6046KJ

q   = \Delta H = -335.6046KJ =- 335604.6J

m = 5692g

t1 = 210C

q = mc\DeltaT

335604.6 = 5692*4.184*(t-21)

t = 35.10C

The final temperature of water = 35.10C >>>.answer